OpenStudy (anonymous):

the square of x is equal to 4 times the square of y. If x is 1 more than twice y , what is the value of x?

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

x^2 = 4*y^2

OpenStudy (anonymous):

(1+2y)^2 = 4y^2

OpenStudy (anonymous):

1 + 2y = 4y

OpenStudy (anonymous):

1/2=y???

OpenStudy (anonymous):

x^2 = 4(1/2)^2

thomaster (thomaster):

no, 1+4y+4y^2=4y^2

OpenStudy (anonymous):

oh so

jimthompson5910 (jim_thompson5910):

x^2 = 4*y^2 x = 1+2y x^2 = 4*y^2 (1+2y)^2 = 4*y^2 (1+2y)^2 = (2y)^2 1+2y = 2y I'll let you finish

OpenStudy (anonymous):

x^2 = 4*y^2 x = 1+2y x^2 = 4*y^2 (1+2y)^2 = 4*y^2 (1+2y)^2 = (2y)^2 1+2y = 2y I'll let you finish

OpenStudy (anonymous):

x^2 = 4*y^2 x = 1+2y x^2 = 4*y^2 (1+2y)^2 = 4*y^2 (1+2y)^2 = (2y)^2 1+2y = 2y I'll let you finish

thomaster (thomaster):

@Michael2392010 Don't copy other peoples answers will ya?

OpenStudy (anonymous):

jk

jimthompson5910 (jim_thompson5910):

whoa now, let's not get racist here...

jimthompson5910 (jim_thompson5910):

Correction, forgot about the absolute values... x^2 = 4*y^2 x = 1+2y x^2 = 4*y^2 (1+2y)^2 = 4*y^2 (1+2y)^2 = (2y)^2 |1+2y| = |2y| 1+2y = 2y or 1+2y = -2y solve both for y

OpenStudy (anonymous):

that doesnt make any since

OpenStudy (anonymous):

u cant

jimthompson5910 (jim_thompson5910):

one of those equations has no solution, but the other does have a solution

OpenStudy (anonymous):

wait question

OpenStudy (anonymous):

how did u get -2y

OpenStudy (anonymous):

oh cuz its sqrt

OpenStudy (anonymous):

nvm i see

jimthompson5910 (jim_thompson5910):

I used the idea that sqrt(x^2) = |x| and if |x| = k, then x = k or x = -k

OpenStudy (anonymous):

1/-4 = y

jimthompson5910 (jim_thompson5910):

or y = -1/4, use this to find x

OpenStudy (anonymous):

x^2 = 4*(-1/4)^2

OpenStudy (anonymous):

x^2 = 4 * (-1/8)

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

x^2 = -4/8

jimthompson5910 (jim_thompson5910):

(-1/4)^2 is NOT -1/8

OpenStudy (anonymous):

oh 16 oops

jimthompson5910 (jim_thompson5910):

it's not 16 either, but that's closer

OpenStudy (anonymous):

x^2 = -1/64

OpenStudy (anonymous):

i ment 16 is the denom

OpenStudy (anonymous):

z thats wrong

jimthompson5910 (jim_thompson5910):

(-1/4)^2 = (-1/4)*(-1/4) the two negatives on the right side will multiply to a positive number

OpenStudy (anonymous):

x^2 = 1/64

OpenStudy (anonymous):

x^2 = 4/64

OpenStudy (anonymous):

x^2 = 1/16

OpenStudy (anonymous):

x = 1/4 :D

OpenStudy (anonymous):

thanks sir

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

remember that x^2 = 4*y^2

jimthompson5910 (jim_thompson5910):

so you should have x^2 = 4*(1/16) x^2 = 1/4

jimthompson5910 (jim_thompson5910):

so x = ??

OpenStudy (anonymous):

1/2?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

you can also use the second equation x = 1+2y x = 1+2(-1/4) x = 1 - 1/2 x = 2/2 - 1/2 x = 1/2